# Defining subsystems in Hilbert spaces with non-Euclidean metric

### v2

@@ Abstract
This work outlines a consistent method of identifying subsystems in finite-dimensional Hilbert spaces, independent of the underlying inner-product structure. It has been well established that Hilbert spaces with modified inner-product, defined through the so-called metric operator, turn out to be the most natural ways to represent certain phenomena such as those involving balanced gain and loss resulting in pseudo-Hermitian Hamiltonians. For composite systems undergoing pseudo-Hermitian evolution, defining the subsystems is generally considered feasible only when the metric operator is chosen to have a tensor product form so that a partial trace operation can be well defined. In this work, we use arguments from algebraic quantum mechanics to show that the subsystems can be well-defined in every metric space – irrespective of whether or not the metric is of tensor product form. This is done by identifying subsystems with a decomposition of the underlying $C^{*}$-algebra into commuting sub-algebras. We show that different subsystem decompositions correspond to choosing different equivalence classes of the GNS representation. Furthermore, given a form of pseudo-Hermitian Hamiltonian, the choice of the Hamiltonian compatible metric characterizes the subsystem decomposition and as a consequence, the entanglement structure in the system. We clarify how each of the subsystems, defined this way, can be tomographically constructed and that these subsystems satisfy the no-signaling principle. With these results, we put all the choices of the metric operator on an equal footing.
@@ I Introduction
It is an axiomatic assumption in quantum theory that the Hilbert space of a composite system has a tensor product decomposition. That is to say, a system with two degrees of freedom, represented by the set of variables $m_{1}$ and $m_{2}$ has the wavefunction $\psi(m_{1},m_{2})\in\mathcal{H}$ in the form $\psi(m_{1},m_{2})=\sum_{i}\psi_{i}(m_{1})\otimes\varphi_{i}(m_{2})\in\mathcal{H}_{1}\otimes\mathcal{H}_{2}$. Such a decomposition allows for a description of the sub-system localized to a sub-space of the total Hilbert space, with operations on this sub-system defined as operations on this sub-space. While this axiom has a footing in some fundamental principles of functional analysis and can be argued from deeper laws of nature  [2, 1], in certain situations, it does not hold good. The most celebrated demonstration of such a deviation is the system of indistinguishable particles [3, 4] and systems with infinite degrees of freedom described by quantum field theory [6, 5]. Another example, and the central topic of this work, is the system that evolves under pseudo-Hermitian Hamiltonians [7, 8]. Pseudo-Hermitian operators are the non-Hermitian operators with real eigenvalues. They arise naturally in certain optical settings with equal gain and loss [9, 10]. It was realized that such evolutions can be seen as unitary evolutions in Hilbert spaces with modified inner product structures induced by the metric operator  [12, 11]. In fact, such a description is necessary to avoid fallacious results like the violation of the no-signaling principle  [13]. For composite systems under such evolutions, the partial trace operation is well-defined if the metric operator allows a tensor product decomposition of the Hilbert space. Such choices of the metric operators are often chosen over other metric operators to define the subsystem  [14, 15]. However, it can so happen that for a given Hamiltonian, none of the choices of the metric operator allow for a tensor product structure [16]. In such situations, claiming that subsystems cannot be defined would be incorrect, especially if the experimentalist can probe the different degrees of freedom independently of each other. Therefore, to define subsystems in such Hilbert spaces one must use an operational definition of the subsystems.
@@
In this work, we outline the problems that arise in defining a subsystem under pseudo-Hermitian evolution and resolve them using tools from the algebraic formulation of quantum mechanics. In the literature, the issue of defining subsystems independently of the Hilbert space structure has been explored by decomposing the underlying $C^{*}$-algebra [17, 18]. We use this algebraic basis to define subsystems in non-trivial inner product spaces. This allows us to put all the choices of the metric operator on equal footing. The central argument of this work is that the different decomposition of the underlying algebra amounts to choosing a particular family of inner product Hilbert spaces to represent a system. Consequently, while a pseudo-Hermitian evolution can be represented as unitary evolution in a family of inner product spaces, different choices of the inner product spaces can result in the probing of different sub-system decompositions. Finally, we show that the no-signaling principle is not broken under this description of the sub-system. While thus far the literature has relied on a genuine tensor product structure of the inner product spaces, we show that under any choice of the metric operator, the no-signaling principle holds for a well-defined sub-system.
@@ I.1 Underlying vector space and the metric
The “pure” state of a classical system at a given time is given by a set of real numbers $\{m_{1},m_{2},...\}$ which correspond to a point in the configuration space. In other words, a measurement of the system can yield these values simultaneously. Let $m_{i}\in M_{i}$ where $M_{i}$ is a set of possible values the state can take. We call a system bipartite if the underlying configuration space is $M=M_{1}\times M_{2}$. This means that the experimentalist has the means to measure the parameters of $M_{1}$ and $M_{2}$ at the same time. For example, by placing the spin detectors at different points in space, one can measure a system’s spin and position. A quantum state is a complex-valued continuous function over the configuration space, given by $\psi(m_{1},m_{2},...)$. The space of all continuous functions on a locally compact Hausdaurff space $M$, denoted by $C^{\infty}(M)$, is a normed vector space that is complete with respect to the metric induced by the norm. This defines a Hilbert space. We are generally interested in a subspace $\mathcal{F}(M)\subset C^{\infty}(M)$, which is also complete with respect to the norm. The elements of this space are the function $f(m_{1},m_{2})$ such that $f(m_{1},\cdot)=f_{2}(\cdot)\in\mathcal{F}_{2}(M_{2})$ and $f(\cdot,m_{2})=f_{1}(\cdot)\in\mathcal{F}_{1}(M_{1})$, where $\mathcal{F}_{1}(M_{1})$ and $\mathcal{F}_{2}(M_{2})$ are set of continuous functions over $M_{1}$ and $M_{2}$. Let the set $\mathcal{F}_{1}(M_{1})\otimes\mathcal{F}_{2}(M_{2})$ denote the closure of the set of linear combinations of functions of the form $\varphi(m_{1},m_{2})=f_{1}(m_{1})\times f_{2}(m_{2})$. Then, we have the equality $\mathcal{F}_{1,2}(M_{1}\times M_{2})=\mathcal{F}_{1}(M_{1})\otimes\mathcal{F}_{2}(M_{2})$ if at least one of the sets $\mathcal{F}_{i}(M_{i})$ is a complete normed vector space   [19].
@@
Therefore, for a system whose underlying configuration space has a Cartesian product form, we assume that the vector space of pure quantum states has a tensor product decomposition. This vector space is converted to a Hilbert space by introducing an inner product map. As we discuss below, the tensor product decomposition of the resulting Hilbert space depends on one’s choice of the inner product map.
I.2 Overview of quantum mechanics in non-Euclidean metric space
@@
Given two elements $\psi$ and $\phi$ on a vector space $\mathcal{V}$, define a Euclidean (or trivial) inner product map $\langle\psi|\phi\rangle\ =\bar{\psi}\phi$, where $\bar{\psi}$ is the conjugate transpose of the element $\psi$. With respect to this inner product, one can define a class of non-Euclidean (or non-trivial) inner products facilitated by a Hermitian, invertible operator $G$ called a metric operator: $\langle\psi|\phi\rangle_{G}\ =\bar{\psi}G\phi=\langle\psi|G\phi\rangle$. We therefore have a class of inner product spaces $\mathcal{H}_{G}:=(\mathcal{V},G)$, where we will denote $\mathcal{H}_{\mathbb{1}}$ simply as $\mathcal{H}$. It can be shown that if $\mathcal{H}_{\mathbb{1}}$ is a Hilbert space, then $\mathcal{H}_{G}$ is also a Hilbert space, i.e. complete with respect to the norm induced by their respective inner product map. The issues in defining subsystems in these Hilbert spaces arise since for an underlying vector space $\mathcal{V}_{1}\otimes\mathcal{V}_{2}$, one has the freedom to endow it with a metric operator that is not in the form of $G_{1}\otimes G_{2}$. In such a case, the partial trace operation on $\mathcal{H}_{G}$ is not well-defined.
An observable is defined as an operator with real expectation values. In $\mathcal{H}_{G}$, such operators must satisfy the relation $G\mathcal{O}=\mathcal{O}^{\dagger}G$, called pseudo-Hermiticity  [8]. An operator can be an observable in two different Hilbert spaces since the metric satisfying the pseudo-Hermiticity condition is not unique. Let $\mathcal{H}_{G^{\prime}}$ be such a Hilbert space. It is characterized by an isometric map $T:\mathcal{H}_{G^{\prime}}\to\mathcal{H}_{G}$ which satisfies the relation $\eta^{\prime}T^{-1}=U\eta$ where $U$ is any unitary map ($U^{\dagger}U=1$). Therefore the isometries $T$ are parameterized by the operator $U$, hence we will denote them by $T_{U}$. The metric operators are related by $T_{U}^{\dagger}GT_{U}=G^{\prime}$, where this relation is independent of $U$. The operator $\mathcal{O}$ is an observable in both metric spaces if and only if $[T_{U},\mathcal{O}]=0$. Given metric operators $G$ and $G^{\prime}$, this relation holds for a unique unitary $U$.
@@
A pseudo-Hermitian operator in $\mathcal{H}_{G}$ can be mapped to a Hermitian operator in $\mathcal{H}$ via $\mathcal{O}\rightarrow\mathcal{O}_{\eta}=\eta\mathcal{O}\eta^{-1}$, where $\eta$ is the unique positive square root of the metric operator $G$. Note that $\eta:\psi\rightarrow\eta\psi$ is an isometry between the spaces $\mathcal{H}_{G}$ and $\mathcal{H}$. Furthermore, since $G^{\prime}$ is another possible metric operator, $\mathcal{O}_{\eta^{\prime}}=\eta^{\prime}\mathcal{O}\eta^{\prime-1}$ is another Hermitisization of the observable $\mathcal{O}$. These two Hermitian forms of $\mathcal{O}$ are unitarily equivalent: $\mathcal{O}_{\eta}=U\mathcal{O}_{\eta^{\prime}}U^{\dagger}$. In fact, one can show that $[T_{U},\mathcal{O}]=0$ if and only if $\mathcal{O}_{\eta}=U\mathcal{O}_{\eta^{\prime}}U^{\dagger}$[16] (see Fig.  1). It can be shown that the space of all allowed metric operators compatible with Hamiltonian $H$, can be parametrized by the set of real vectors $\vec{\lambda}$[11]. Given $\{\psi_{i}\}$ and $\{\phi_{j}\}$ the left and right eigenvectors of $H$ respectively, (given it has a non-degenerate spectrum), the most general metric is given by $G_{\lambda}=\sum_{i}\lambda_{i}|\psi_{i}\rangle\langle\phi_{i}|$.
Figure 1: Isometric maps between different metric Hilbert spaces. The state $\psi$ in $\mathcal{H}_{G^{\prime}}$ can be mapped to a state in $\mathcal{H}$ using isometric maps $T_{U}$. The resulting states are related by a unitary $U$ in $\mathcal{H}$.
I.3 Open-system perspective
@@
Pseudo-Hermitian evolutions in quantum systems have been experimentally exhibited many times in recent years  [20, 21, 22]. This is typically achieved by introducing an interacting potential wherein the amplitude and frequency of the gain and loss of the probability current are carefully controlled to maintain the potential’s PT-symmetry  [23]. That is, the interacting Hamiltonian is invariant under the action of the parity and time reversal operations. Within certain parameters (referred to as the ”unbroken” PT-symmetry regime) the Hamiltonian has a Pseudo-Hermitian form  [24]. Due to the presence of interaction terms in the Hamiltonian, such an evolution is sometimes seen as open system dynamics, even though the effective Hamiltonian has real eigenvalues. A quintessential feature that separates a pseudo-Hermitian evolution from a genuinely non-Hermitian (non-unitary) evolution is the revival of complete information of the system throughout the dynamics, measured as probability values or the information back-flow. It is argued  [25, 26] that such a complete revival is due to the presence of a finite-sized entangled partner in the environment of the system, where the Hermitisization operation $\psi\rightarrow\eta\psi$ is seen as an embedding of the system into a larger Hermitian system. This description, while illuminating, views pseudo-Hermiticity as a non-trivial extension of quantum mechanics. As we will elucidate in the subsequent sections, the pseudo-Hermitian Hamiltonian, although simulated through interaction terms, induces a unitary evolution.
A large class of open system dynamics follows a Lindblad-type master equation, which can be recast as an evolution under an effective non-Hermitian Hamiltonian and a “jump-term”. Post-selecting on the trajectories without a jump, the evolution of the system can be cast as an evolution under the non-Hermitian Hamiltonian. If the effective Hamiltonian is pseudo-Hermitian, then one can show that the Lindblad equation is simply the von Neumann equation for a redefined pseudo-Hermitian state, and the jump term vanishes (see A for details), effectively reproducing the unitary dynamics.

### v3

# Decomposition of a system in pseudo-Hermitian quantum mechanics

@@ Abstract
This work outlines a consistent method of identifying subsystems in finite-dimensional Hilbert spaces, independent of the underlying inner-product structure. Such Hilbert spaces arise in $\mathcal{P}\mathcal{T}$-symmetric quantum mechanics, where a non-Hermitian Hamiltonian is made self-adjoint by changing the inner product using the so-called “metric operator”. This is the framework of pseudo-Hermitian quantum mechanics. For composite quantum systems in this framework, defining subsystems is generally considered feasible only when the metric operator is chosen to have a tensor product form so that a partial trace operation can be well defined. In this work, we use arguments from algebraic quantum mechanics to show that the subsystems can be well-defined in every metric space – irrespective of whether or not the metric is of tensor product form. This is done by identifying subsystems with a decomposition of the underlying $C^{*}$-algebra into commuting subalgebras. Although the choice of the metric is known to have no effect on the system’s statistics, we show that different choices of the metric can lead to inequivalent subsystem decompositions. Each of the subsystems can be tomographically constructed and these subsystems satisfy the no-signalling principle. With these results, we put all the choices of the metric operator on an equal footing for composite systems.
@@ I Introduction
It is an axiomatic assumption in quantum theory that the Hilbert space of a composite system has a tensor product decomposition. That is to say, a system with two degrees of freedom, represented by the set of variables $m_{1}$ and $m_{2}$ has the wavefunction $\psi(m_{1},m_{2})\in\mathcal{H}$ in the form $\psi(m_{1},m_{2})=\sum_{i}\psi_{i}(m_{1})\otimes\varphi_{i}(m_{2})\in\mathcal{H}_{1}\otimes\mathcal{H}_{2}$. Such a decomposition allows for a description of the subsystem localized to a subspace of the total Hilbert space, with operations on this subsystem defined as operations on this subspace. While this axiom has a footing in some fundamental principles of functional analysis and can be argued from deeper laws of nature  [2, 1], in certain situations, it does not hold good. The most celebrated demonstration of such a deviation is the system of indistinguishable particles [3, 4] and systems with infinite degrees of freedom described by quantum field theory [6, 5]. Another example, and the central topic of this work, is the system that evolves under $\mathcal{P}\mathcal{T}$-symmetric or pseudo-Hermitian Hamiltonians [7, 8]. Pseudo-Hermitian operators are the non-Hermitian operators with eigenvalues in complex conjugate pairs or real eigenvalues (quasi-Hermitian). They arise naturally in certain optical settings with equal gain and loss [9, 10]. It was realized that such evolutions can be seen as unitary evolutions in Hilbert spaces with modified inner product structures induced by a metric operator[28, 12]. In fact, such a description is necessary to avoid fallacious results like the violation of the no-signalling principle  [13].
For composite systems under such evolutions, the partial trace operation is well-defined if the metric operator allows a tensor product decomposition of the Hilbert space. Such choices of the metric operators are often chosen over other metric operators to define the subsystem  [14, 15]. However, it can so happen that for a given Hamiltonian, none of the choices of the metric operator allow for a tensor product structure [16]. In such situations, claiming that subsystems cannot be defined would be incorrect, especially if the experimentalist can probe the different degrees of freedom independently of each other. Therefore, to define subsystems in such Hilbert spaces one must use an operational definition of the subsystems, which is the main goal of this article.
@@
In this work, we outline the problems that arise in defining a subsystem in pseudo-Hermitian quantum mechanics and resolve them using tools from the algebraic formulation of quantum mechanics. In the literature, the issue of defining subsystems independently of the Hilbert space structure has been explored by decomposing the underlying $C^{*}$-algebra [17, 18]. We use this algebraic basis to define subsystems in non-Euclidean inner product spaces. This allows us to put all the choices of the metric operator on equal footing. The central argument of this work is that different decompositions of the underlying algebra correspond to particular families of inner product spaces used to represent the system. Consequently, while a quasi-Hermitian evolution can be represented as unitary evolution in a family of inner product spaces, different choices of the inner product spaces can result in the probing of different subsystem decompositions. While thus far the literature has relied on a genuine tensor product structure of the inner product spaces, we show that under any choice of the metric operator, the no-signalling principle holds for a well-defined subsystem.
The article is structured as follows: in this section, we will review the aspects of pseudo-Hermitian quantum mechanics. In section II we will give a short introduction to $C^{*}$-algebras and their pseudo-Hermitian representations. We will define bi-partitions of a system, in a given inner product space and show the dependence of the partitioning on the choice of the inner product in section III. In section IV we will show that local tomography can be performed for systems in any metric space, and the no-signalling principle is satisfied. We conclude with a short discussion in section V.
@@ I.1 Underlying vector space and the metric
The “pure” state of a classical system at a given time is given by a set of real numbers $\{m_{1},m_{2},...\}$ that correspond to a point in the configuration space. In other words, a measurement of the system can yield these values simultaneously. Let $m_{i}\in M_{i}$ where $M_{i}$ is a set of possible values the state can take. We call a system bipartite if the underlying configuration space is $M=M_{1}\times M_{2}$. This means that the experimentalist has the means to measure the parameters of $M_{1}$ and $M_{2}$ at the same time. For example, by placing the spin detectors at different points in space, one can measure the spin and position of a system. A quantum state is a complex-valued continuous function over the configuration space, given by $\psi(m_{1},m_{2},...)$. If the configuration space is a compact subset of $\mathbb{R}^{n}$ (or more generally a locally compact Hausdorff space) then the space of smooth functions is a vector space $C^{\infty}(M)$, where one can also define a “norm” for the vectors. This defines a Hilbert space. We are generally interested in a subspace $\mathcal{F}(M)\subset C^{\infty}(M)$. The elements of this space are the function $f(m_{1},m_{2})$ such that $f(m_{1},\cdot)=f_{2}(\cdot)\in\mathcal{F}_{2}(M_{2})$ and $f(\cdot,m_{2})=f_{1}(\cdot)\in\mathcal{F}_{1}(M_{1})$, where $\mathcal{F}_{1}(M_{1})$ and $\mathcal{F}_{2}(M_{2})$ are set of continuous functions over $M_{1}$ and $M_{2}$. Let the set $\mathcal{F}_{1}(M_{1})\otimes\mathcal{F}_{2}(M_{2})$ denote the closure of the set of linear combinations of functions of the form $\varphi(m_{1},m_{2})=f_{1}(m_{1})\times f_{2}(m_{2})$. Then, we have the equality $\mathcal{F}_{1,2}(M_{1}\times M_{2})\cong\mathcal{F}_{1}(M_{1})\otimes\mathcal{F}_{2}(M_{2})$ if at least one of the sets $\mathcal{F}_{i}(M_{i})$ is a complete normed vector space   [19].
@@
These results from functional analysis lead us to conclude that for a system whose underlying configuration space has a Cartesian product form, the vector space of pure quantum states should have a tensor product decomposition. In this way, one motivates the tensor product postulate for composite quantum systems. This vector space is converted to a Hilbert space by introducing an inner product map. As we discuss below, the tensor product decomposition of the resulting Hilbert space depends on one’s choice of the inner product map.
I.2 Overview of pseudo-Hermitian quantum mechanics
@@
Given two elements $\psi$ and $\phi$ in a vector space $\mathcal{V}$, we define a Euclidean inner product map $\langle\psi|\phi\rangle\ =\psi^{\dagger}\phi$, where $\psi^{\dagger}$ is the conjugate transpose of the element $\psi$. With respect to this inner product, one can define a class of non-Euclidean inner products facilitated by a Hermitian, positive-definite operator $G$ called a metric operator: $\langle\psi|\phi\rangle_{G}\ :=\psi^{\dagger}G\phi=\langle\psi|G\phi\rangle$. This gives us a class of inner product spaces $\mathcal{H}_{G}:=(\mathcal{V},G)$, where we will denote $\mathcal{H}_{\mathbb{1}}$ simply as $\mathcal{H}$. It can be shown that if $\mathcal{H}_{\mathbb{1}}$ is a Hilbert space, then $\mathcal{H}_{G}$ is also a Hilbert space, i.e. complete with respect to the norm induced by their respective inner product map. If the underlying vector space is of the form $\mathcal{V}_{1}\otimes\mathcal{V}_{2}$, then the Hilbert space $\mathcal{H}_{G}=(\mathcal{V}_{1}\otimes\mathcal{V}_{2},G)$ has a tensor product decomposition $\mathcal{H}_{G}=\mathcal{H}_{G_{1}}\otimes\mathcal{H}_{G_{2}}$ if and only if the metric operator has he form $G=G_{1}\otimes G_{2}$ where $G_{1}$ is metric operator on $\mathcal{V}_{1}$ and similarly for $G_{2}$. Given a density matrix on $\mathcal{H}_{G}$, its subsystem is usually characterized by a completely positive and trace-preserving (CPTP) map to a smaller Hilbert subspace through the partial trace operation. However, such a map is well defined only if the Hilbert space $\mathcal{H}_{G}$ has a tensor product decomposition of the form $\mathcal{H}_{G_{1}}\otimes\mathcal{H}_{G_{2}}$. For Hilbert spaces with a metric operator not in the tensor product form, the partial trace operation is not well defined. In this work, we will outline a method to construct the subsystem statistics for the systems defined on such Hilbert spaces.
Hilbert spaces with non-Euclidean metric operators arise naturally in $\mathcal{P}\mathcal{T}$-symmetric quantum mechanics. For example, the evolution of systems under balanced gain and loss can be seen as generated by a non-Hermitian Hamiltonian. Such Hamiltonians are symmetric under the consecutive action of parity and time reversal operations, and can have real eigenvalues. All such Hamiltonians $H$ satisfy a condition called $G$-pseudo-Hermiticity given by $H^{\dagger}G=GH$ for a Hermitian operator $G$. If, for a given Hamiltonian $H$, there exists a positive-definite operator $G$ satisfying the pseudo-Hermiticity condition, then $H$ is guaranteed to have real eigenvalues []. Such operators are called quasi-Hermitian operators. It turns out that quasi-Hermitian operators have real expectation values in Hilbert spaces with metric operator $G$, and therefore can be seen as observables in $\mathcal{H}_{G}$ as we show below:
$\langle\psi|H\psi\rangle_{G}$ $=\langle\psi|GH\psi\rangle$ (1)
$=\langle\psi|H^{\dagger}G\psi\rangle$
$=\langle H\psi|G\psi\rangle$
$=\langle H\psi|\psi\rangle_{G}$
In pseudo-Hermitian quantum mechanics, the physical Hilbert space is a metric space with a metric operator $G\neq\mathbb{1}$. In such a space a $G$-pseudo-Hermitian operator is also a quasi-Hermitian operator and can be treated as an observable.
@@
An operator can be an observable in two different Hilbert spaces since the metric operator satisfying the pseudo-Hermiticity condition for a given operator is not unique [16]. Let $\mathcal{H}_{G^{\prime}}$ and $\mathcal{H}_{G}$ be two such Hilbert spaces for an observable $\mathcal{O}$, i.e. $\mathcal{O}^{\dagger}G=G\mathcal{O}$ , and $\mathcal{O}^{\dagger}G^{\prime}=G^{\prime}\mathcal{O}$. Then, there exists a linear operator $T$ such that $T^{\dagger}GT=G^{\prime}$ and $[T,\mathcal{O}]=0$. It can be shown that the space of all allowed metric operators compatible with Hamiltonian $H$, can be parametrized by the set of real vectors $\vec{\lambda}$[11]. Given $\{\psi_{i}\}$ and $\{\phi_{j}\}$ the left and right eigenvectors of $H$ respectively, (given it has a non-degenerate spectrum), the most general metric is given by $G_{\lambda}=\sum_{i}\lambda_{i}|\psi_{i}\rangle\langle\phi_{i}|$.
A pseudo-Hermitian operator in $\mathcal{H}_{G}$ can be mapped to a Hermitian operator in $\mathcal{H}$ via the Hermitisization map
$\mathcal{O}\rightarrow\mathcal{O}_{\eta}=\eta\mathcal{O}\eta^{-1},$ (2)
where $\eta$ is the unique positive square root of the metric operator $G$. Note that $\eta:\psi\rightarrow\eta\psi$ is an isometry between the spaces $\mathcal{H}_{G}$ and $\mathcal{H}$. Furthermore, since $G^{\prime}$ is another possible metric operator, $\mathcal{O}_{\eta^{\prime}}=\eta^{\prime}\mathcal{O}\eta^{\prime-1}$ is another Hermitisization of the observable $\mathcal{O}$.
For quasi-Hermitian Hamiltonians, Mostafazadeh’s pseudo-Hermitian quantum mechanics [11] seems to be the most natural formalism. However, general pseudo-Hermitian Hamiltonians with complex eigenvalues leave room for a wider choice of inner product spaces and have been applied to various physical systems [20, 21, 22, 23]
I.3 Open-system perspective on pseudo-Hermitian Hamiltonians
@@
Pseudo-Hermitian evolutions in quantum systems have been experimentally exhibited many times in recent years  [24, 25, 26]. This is typically achieved by introducing an interacting potential wherein the amplitude and frequency of the gain and loss of the probability current are carefully controlled to maintain the potential’s $\mathcal{P}\mathcal{T}$-symmetry  [27]. That is, the interacting Hamiltonian is invariant under the action of the parity and time reversal operations. This is equivalent to maintaining a balanced gain and loss of amplitudes. Within certain parameters (referred to as the “unbroken” $\mathcal{P}\mathcal{T}$-symmetry regime) the Hamiltonian has a quasi-Hermitian form  [28]. Due to the presence of interaction terms in the Hamiltonian, such an evolution is sometimes seen as an open system dynamics, even though the effective Hamiltonian has real eigenvalues. A quintessential feature that separates a quasi-Hermitian evolution from a genuinely non-Hermitian (non-unitary) evolution is the revival of complete information of the system throughout the dynamics, measured as probability values or the information back-flow. It has been argued  [29, 30] that such a complete revival is due to the presence of a finite-sized entangled partner in the environment of the system, where the Hermitisization operation $\psi\rightarrow\eta\psi$ is seen as an embedding of the system in a larger Hermitian system. As we will elucidate in the subsequent sections, the quasi-Hermitian Hamiltonian, although simulated through interaction terms, induces a unitary evolution.
A large class of open system dynamics follows a Lindblad-type master equation, which can be recast as an evolution under an effective non-Hermitian Hamiltonian and a “jump-term”. Post-selecting on the trajectories without a jump, the evolution of the system can be cast as an evolution under the non-Hermitian Hamiltonian. If the effective Hamiltonian is quasi-Hermitian, then one can show that the Lindblad equation is simply the von Neumann equation for a redefined quasi-Hermitian state, and the jump term vanishes (see A for details), effectively reproducing the unitary dynamics. On the other hand, in the “broken regime” where the Hamiltonian has complex eigenvalues, interpreting the Hamiltonian as a generator of time translation can be tricky [31]. The compatible operator $G$ is not positive definite in this regime, and as a result the corresponding inner product space is not a well-defined Hilbert space. One can also induce a time-dependent metric operator e.g. [32], or use the Euclidean inner product [33, 34, 35] to study such Hamiltonians. In this work we will restrict ourselves to the formalism of pseudo-Hermitian quantum mechanics, where the norm is induced by a non-Euclidean positive-definite inner product. Therefore, the form of the Hamiltonian or the regimes of broken or unbroken phases is not of particular importance to us. However, our results can help understand the properties of the composite systems evolving under such Hamiltonians if treated under the pseudo-Hermitian formalism.